Circular No. 8175

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003gv IN MCG +05-3-66 D. Weisz and W. Li report the LOTOSS/KAIT (cf. IAUC 8172) dis-

covery of an apparent supernova (mag ∼ 18.4), on images from Aug. 6.5 and 7.5 UT, located at α = 1^{h}08^{m}53^{s}.18, δ = +32^{o}05′59′′.3 (equinox 2000.0), which is 11′′.1 east and 0′′.3 south of the nucleus of MCG +05-3-66. A KAIT image from Jan. 24.2 showed nothing at this position (limiting mag ∼ 19.5).

SUPERNOVAE 2003gt AND 2003gu A. V. Filippenko, R. J. Foley, and L. Desroches, University of Califor-

nia, Berkeley, report that inspection of CCD spectra (range 320–1000 nm), obtained on Aug. 5 UT with the Shane 3-m telescope at Lick Observatory, reveals that SN 2003gt (IAUC 8172) is of type Ia, close to maximum brightness. The expansion velocity measured from the Si II 635.5-nm absorption line is ∼ 11000 km/s. SN 2003gu (IAUC 8173) appears to be of type II, shortly after the explosion; the continuum is blue and generally featureless, and Hα exhibits a P-Cyg profile. However, the He I 587.6-nm absorption is unusually strong; thus, the spectrum somewhat resembles that of a young type-IIb supernova. The expansion velocities measured from the Hα and He I absorption lines are ∼ 17000 and 10000 km/s, respectively.

SUPERNOVA 2003gs IN NGC 936 J. Mikolajewska and A. Szostek, Copernicus Astronomical Center,

Warsaw, report preliminary near-infrared magnitudes for SN 2003gs, obtained with the 1.55-m Carlos Sánchez Telescope (+ CAIN II) at Tenerife: Aug. 1.24 UT: J = 13.0 ± 0.1; Aug. 2.22, J = 13.5, K_{s} = 13.1; Aug. 3.21, J = 13.6, K_{s} = 13.1; Aug. 4.21, J = 13.6, H = 13.0, K_{s} = 13.0.

V2573 OPHIUCHI D. K. Lynch, R. J. Rudy, S. Mazuk, and C. C. Venturini, The Aero-

space Corporation; R. B. Perry, Langley Research Center, NASA; and R. C. Puetter, University of California at San Diego, report 0.9–2.5-μm spectroscopy of V2573 Oph from the Lick 3.05-m telescope (+ Aerospace Near Infrared Imaging Spectrograph) on July 26.25 UT. The nova’s spectrum showed emission lines of H I, weak He I (1.0830 and 2.0581 μm), strong O I and N I, and moderate Fe II. The first overtone of CO at 2.3 μm was seen weakly in emission. No P-Cyg profiles were detected in any line. The FWHM of the lines was 1400 ± 100 km/s.

2003 August 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green